OpenStudy (anonymous):

x + (x + 8)(−8) = x − 8

OpenStudy (anonymous):

x+-8x-64=x-8

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

-8x-56=0 y=-7

OpenStudy (anonymous):

x-8x-64=x-8 -8x-64+8=0 -8x+56=0 x=-7

OpenStudy (anonymous):

tsk tsk raheen

OpenStudy (anonymous):

:/

hero (hero):

That's the first mistake computer ever made